OpenStudy (anonymous):

What is the equation of the line passing through point (3, -1) and parallel to the given line? 3x – y = 5

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@texaschic101

OpenStudy (igreen):

First, re-arrange it into slope-intercept form. 3x - y = 5 Subtract 3x to both sides: -y = -3x + 5 Multiply -1 to both sides: y = 3x - 5 What's the slope here?

OpenStudy (anonymous):

y=3x +6

OpenStudy (anonymous):

nvm

OpenStudy (igreen):

No, the slope is represented by one number. y = 3x - 5 This is in slope-intercept form, y = mx + b, where 'm' is the slope. So what's the slope here?

OpenStudy (anonymous):

y=3x+ 5?

OpenStudy (anonymous):

3,5?

OpenStudy (igreen):

No.. The slope is right before 'x'

OpenStudy (igreen):